OpenStudy (anonymous):

Find an equation of the circle that satisfies the given conditions. Endpoints of a diameter are P(−2, 1) and Q(8, 9).

OpenStudy (zzr0ck3r):

equation of a circle\[(x-x_0)^2+(y-y_0)^2=r^2\]

OpenStudy (zzr0ck3r):

now use the points given to find the radius

OpenStudy (anonymous):

the circle has that point as center which is the midpoint of diameter and radius is distance between diameter ends/2

OpenStudy (anonymous):

what?

OpenStudy (zzr0ck3r):

the equation of a circle is \[(x-x_0)^2+(y-y_0)^2=r^2\\where\space (x_0,y_0)\text{ is the center}\\\text{to find the radius we use the distance formula with the two points given}\] \[r=\sqrt{(8-(-2))^2+(9-1)^2}=\sqrt{164}\]so you have \[(x-8)^2+(y-9)^2=164\]

OpenStudy (anonymous):

so that is the answer?